OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6).

OpenStudy (texaschic101):

y = -3x + 12 In y = mx + b form, the m is the slope. In this equation, -3 is the slope. A parallel line will have the same slope, so the slope we will use is -3. y = mx + b (-1,6) x = -1 and y = 6 slope(m) = -3 now we sub 6 = -3(-1) + b 6 = 3 + b 6 - 3 = b 3 = b your parallel equation is : y = -3x + 3 or if you need it in standard form, it is : 3x + y = 3

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

sure thing :)